OpenStudy (anonymous):

In triangle ABC, measure A = X, measure B = 2X+2, and measure C = 3X+4. What is the value of X?

jimthompson5910 (jim_thompson5910):

For any triangle, all three angles always add to 180 degrees.

jimthompson5910 (jim_thompson5910):

A+B+C = 180 x + (2x+2) + (3x+4) = 180 x + 2x + 2 + 3x + 4 = 180 6x + 6 = 180 x = ??